### v1

@@ B.2 Factor estimation asympototic properties
Under Assumption A.2 and A.3, factor estimation error centered against the normalized true factor converges to a normal distribution at the rate of $\sqrt{N}$: as $N,T\to\infty$ for $\forall t$,
@@
$\sqrt{N}\left(\bm{\hat{f}}_{t}-\bm{f}_{t}\right)$ $=\sqrt{N}\left(\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\left(X_{t}\hat{\Gamma}\bm{f}_{t}+\bm{\tilde{\epsilon}}_{t}\right)-\bm{f}_{t}\right)$
$=\sqrt{N}\left(\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\left(X_{t}\hat{\Gamma}\bm{f}_{t}\right)-\bm{f}_{t}\right)+\sqrt{N}\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\bm{\tilde{\epsilon}}_{t}$
where $\bm{\tilde{\epsilon}}_{t}$ is the estimated error term with estimated $\hat{\Gamma}$ and true $\bm{f}_{t}$. Given Proposition 1, $\hat{\Gamma}-\hat{\Gamma}^{0}=\mathcal{O}_{p}\left(1/\sqrt{NT}\right)$. The first term is simply $\mathcal{O}_{p}\left(1/\sqrt{NT}\right)$. For the second term:
$\sqrt{N}\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\bm{\tilde{\epsilon}}_{t}=$ $\sqrt{N}\left(\Gamma^{\prime}X^{\prime}_{t}X_{t}\Gamma\right)^{-1}\Gamma^{\prime}X_{t}^{\prime}\bm{\epsilon_{t}}+\mathcal{O}_{p}(1)$
@@
Assumption A.3 delineates the properties of $\Omega_{t}^{xx}$ and $\Omega_{t}^{x\epsilon}$. Notably, despite the presence of multiple matrix multiplications, the matrices $\Omega_{t}^{xx}$ and $\Omega_{t}^{x\epsilon}$ remain invariant under these operations. Consequently, the term $\mathbb{V}_{t}^{[2]}$ is constant across all observational units.
@@ B.3 Consistency and asymptotic property of the ATT estimation
$=\textbf{x}_{it}\Gamma\bm{f}^{\prime}_{t}-\textbf{x}_{it}\hat{\Gamma}\hat{\bm{f}}^{\prime}_{t}+\epsilon_{it},$
$=\textbf{x}_{it}\left(\left(\mathcal{I}_{L}\otimes\bm{f}_{t}\right)\bm{\gamma}-(\mathcal{I}_{L}\otimes\hat{\bm{f}}_{t})\hat{\bm{\gamma}}\right)+\epsilon_{it},$
$=\textbf{x}_{it}\left(\left(\mathcal{I}_{L}\otimes\bm{f}_{t}\right)\bm{\gamma}-\mathcal{I}_{L}\otimes(\bm{f}_{t}+\bm{e}_{f_{t}})(\bm{\gamma}+\textbf{e}_{\gamma})\right)+\epsilon_{it},$
$=\textbf{x}_{it}\left((\mathcal{I}_{L}\otimes\bm{f}_{t})\textbf{e}_{\gamma}-(\mathcal{I}_{L}\otimes\bm{e}_{f_{t}}\bm{\gamma})-(\mathcal{I}_{L}\otimes\bm{e}_{f_{t}})\textbf{e}_{\gamma}\right)+\epsilon_{it}$
$=\textbf{x}_{it}E_{\Gamma}\bm{f}^{\prime}_{t}-\textbf{x}_{it}\Gamma\bm{e}^{\prime}_{f_{t}}-\textbf{x}_{it}E_{\Gamma}\bm{e}^{\prime}_{f_{t}}+\epsilon_{it},$
@@
where $\bm{e}_{f_{t}}=\bm{f}_{t}-\hat{\bm{f}}_{t}$ is a vector of estimation error of the factor $\bm{f}_{t}$, $\bm{e}_{\gamma}$ is vectorized estimation error of the mapping matrix $E_{\Gamma}=\Gamma-\hat{\Gamma}$. The third step converts the vector-matrix multiplication into vector multiplications with the Kronecker product, $\bm{x}_{it}\Gamma\bm{f}^{\prime}_{t}=\bm{x}_{it}(\mathcal{I}_{L}\otimes\bm{f}_{t})\bm{\gamma}$, where $\mathcal{I}_{L}$ is an $L\times L$ identity matrix. The bias of the estimated ATT is the sum of four terms $A_{1,it}$, $A_{2,it}$, $A_{3,it}$, and $\epsilon_{it}$. By proposition 1 and 2, we have the following results:
$A_{1,it}=\bm{x}_{it}E_{\Gamma}\bm{f}^{\prime}_{t}=\mathcal{O}_{p}\left(1/\sqrt{N_{treat}T_{pre}}\right).$
$A_{2,it}=-\bm{x}_{it}\Gamma\bm{e}^{\prime}_{f_{t}}=\mathcal{O}_{p}\left(1/\sqrt{N_{ctrl}}\right).$
$A_{3,it}=-\bm{x}_{it}E_{\Gamma}\bm{e}^{\prime}_{f_{t}}=\mathcal{O}_{p}\left(1/\sqrt{N_{treat}T_{pre}}\right).$
Since we estimate the factor $\bm{f}_{t}$ using only control units and update the mapping matrix $\Gamma$ with treated units in the pre-treatment period, both $\bm{f}_{t}$ and $\Gamma$ converge over different dimensions of $T$ and $N$. Consequently, the error term $\epsilon_{it}$ is assumed to have zero means, leading to the bias of the estimated ATT also converging to zero:
$\hat{\delta}_{it}-\delta_{it}$ $=\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{treat}T_{pre}}}\right)+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{ctrl}}}\right)+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{treat}T_{pre}}}\right)+\epsilon_{it}$
@@ C.2 Finite sample properties
In this section, we provide additional simulation results to investigate the finite sample properties of the CSC-IFE and SCM methods for comparison.
@@
Appendix D Empirical Application

### v2

@@ B.2 Factor estimation asympototic properties
Under Assumption 2 and 3, factor estimation error centered against the normalized true factor converges to a normal distribution at the rate of $\sqrt{N}$: as $N,T\to\infty$ for $\forall t$,
@@
$\sqrt{N}\left(\hat{F}_{t}-F_{t}\right)$ $=\sqrt{N}\left(\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\left(X_{t}\hat{\Gamma}F_{t}+\tilde{\epsilon}_{t}\right)-F_{t}\right)$
$=\sqrt{N}\left(\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\left(X_{t}\hat{\Gamma}F_{t}\right)-F_{t}\right)+\sqrt{N}\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\tilde{\epsilon}_{t}$
where $\tilde{\epsilon}_{t}$ is the estimation error term with estimated $\hat{\Gamma}$ and true $F_{t}$. Given Proposition 1, $\hat{\Gamma}-\hat{\Gamma}^{0}=\mathcal{O}_{p}\left(1/\sqrt{NT}\right)$. The first term is simply $\mathcal{O}_{p}\left(1/\sqrt{NT}\right)$. For the second term:
$\sqrt{N}\left(\hat{\Gamma}^{\prime}X^{\prime}_{t}X_{t}\hat{\Gamma}\right)^{-1}\hat{\Gamma}^{\prime}X_{t}^{\prime}\tilde{\epsilon}_{t}=$ $\sqrt{N}\left(\Gamma^{\prime}X^{\prime}_{t}X_{t}\Gamma\right)^{-1}\Gamma^{\prime}X_{t}^{\prime}\epsilon_{t}+\mathcal{O}_{p}(1)$
@@
Assumption 3 delineates the properties of $\Omega_{t}^{xx}$ and $\Omega_{t}^{x\epsilon}$. Notably, despite the presence of multiple matrix multiplications, the matrices $\Omega_{t}^{xx}$ and $\Omega_{t}^{x\epsilon}$ remain invariant under these operations. Consequently, the term $\mathbb{V}_{t}^{[2]}$ is constant across all observational units.
@@ B.3 Consistency and asymptotic property of the ATT estimation
$=X_{it}\Gamma F^{\prime}_{t}-X_{it}\hat{\Gamma}\hat{F}^{\prime}_{t}+\epsilon_{it},$
$=X_{it}\left(\left(\mathcal{I}_{L}\otimes F_{t}\right)\bm{\gamma}-(\mathcal{I}_{L}\otimes\hat{F}_{t})\hat{\bm{\gamma}}\right)+\epsilon_{it},$
$=X_{it}\left(\left(\mathcal{I}_{L}\otimes F_{t}\right)\bm{\gamma}-\mathcal{I}_{L}\otimes(F_{t}+e_{f_{t}})(\bm{\gamma}+e_{\gamma})\right)+\epsilon_{it},$
$=-X_{it}\left((\mathcal{I}_{L}\otimes F_{t})e_{\gamma}-(\mathcal{I}_{L}\otimes e_{f_{t}}\bm{\gamma})-(\mathcal{I}_{L}\otimes e_{f_{t}})e_{\gamma}\right)+\epsilon_{it}$
$=-X_{it}e_{\Gamma}F^{\prime}_{t}-X_{it}\Gamma e^{\prime}_{f_{t}}-X_{it}e_{\Gamma}e^{\prime}_{f_{t}}+\epsilon_{it},$
@@
where $e_{f_{t}}=F_{t}-\hat{F}_{t}$ is a vector of estimation error of the factor $F_{t}$, $e_{\gamma}$ is vectorized estimation error of the mapping matrix $e_{\Gamma}=\Gamma-\hat{\Gamma}$. The third step converts the vector-matrix multiplication into vector multiplications with the Kronecker product, $X_{it}\Gamma F^{\prime}_{t}=X_{it}(\mathcal{I}_{L}\otimes F_{t})\bm{\gamma}$, where $\mathcal{I}_{L}$ is an $L\times L$ identity matrix. The bias of the estimated ATT is the sum of four terms $A_{1,it}$, $A_{2,it}$, $A_{3,it}$, and $\epsilon_{it}$. By proposition 1 and 2, we have the following results:
$A_{1,it}=-X_{it}E_{\Gamma}F^{\prime}_{t}=\mathcal{O}_{p}\left(1/\sqrt{N_{treat}T_{pre}}\right).$
$A_{2,it}=-X_{it}\Gamma e^{\prime}_{f_{t}}=\mathcal{O}_{p}\left(1/\sqrt{N_{ctrl}}\right).$
$A_{3,it}=-X_{it}E_{\Gamma}e^{\prime}_{f_{t}}=\mathcal{O}_{p}\left(1/\sqrt{N_{treat}T_{pre}N_{ctrl}}\right).$
Since we estimate the factor $F_{t}$ using only control units and update the mapping matrix $\Gamma$ with treated units in the pre-treatment period, both $F_{t}$ and $\Gamma$ converge over different dimensions of $T$ and $N$. Consequently, the error term $\epsilon_{it}$ is assumed to have zero means, leading to the bias of the estimated ATT also converging to zero:
$\hat{\delta}_{it}-\delta_{it}$ $=\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{treat}T_{pre}}}\right)+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{ctrl}}}\right)+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N_{treat}T_{pre}N_{ctrl}}}\right)+\epsilon_{it}$
@@ C.2 Finite sample properties
In this section, we present additional simulation results to investigate the finite sample properties of the CSC-IFE and SCM methods for comparison. As the results show, the SCM method exhibits bias due to the specific settings of our data generating processes. While the CSC-IFE method performs comparably to the CSC-IPCA method when all predictive covariates are observed, its bias grows significantly as the number of unobserved covariates increases, leading to increasingly unreliable estimates.
@@
Appendix D Empirical Applicatoin